Order posts by limited to posts

12 Mar 2015 12:01:00
12 Mar 2015 09:50:23
TalkTalk have a a fibre outage affecting EoFTTC lines in the East Midlands and East Anglia area. Lines will be in sync, but will not have any further connectivity. TalkTalk are aware and have engineers investigating. They have their own status post here: http://solutions.opal.co.uk/network-status-report.php?reportid=6291
12 Mar 2015 10:18:51
Not further updates as yet.
12 Mar 2015 10:20:37
Update from TT: Our NOC and Network Support Engineers have confirmed that corrective work to restore services to all impacted exchanges remains on-going.
12 Mar 2015 11:44:30
Update from TalkTalk: "Our NOC Engineers have confirmed that all network alarms have cleared and normal working service is now fully restored to all impacted exchanges and customers. There are continued reports of some B2B circuits continuing to experience issues and our IP Engineers are currently investigating these reports. Incident Management will continue to liaise with our NOC Engineers and further updates will be provided upon receipt."
12 Mar 2015 14:20:48
Service has been restored.
Started 12 Mar 2015 00:10:00
Closed 12 Mar 2015 12:01:00

24 Feb 2014 16:00:00
24 Feb 2014 16:01:15
Some work was done today (following issues at the weekend). This work is on part of a redundant pair of routers, and so should have no impact. The normal redundancy aspects have worked, with changes to routing and VRRP, but for some unknown reason there are issues with BGP announcements internally which are causing blips on external connectivity to some Ethernet customers. The work has been completed but we are now investigating how this has had an impact on services so they can be avoided in future.
Started 24 Feb 2014 15:45:00
Closed 24 Feb 2014 16:00:00
Previously expected 24 Feb 2014 16:00:00

02 Sep 2011 09:41:55
02 Sep 2011 09:43:14

The usage graphs were disrupted yesterday due to various work on core routers, and will be reset today starting shortly. This is to fix a problem where the ping monitoring was not working correctly. Service should otherwise be unaffected by this change.

Started 02 Sep 2011
Closed 02 Sep 2011 09:41:55